40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 20     Entry time: Fri Oct 26 21:48:40 2007
Author: waldman 
Type: Configuration 
Category: OMC 
Subject: Fiber to 056 
I set up a 700 mW NPRO in Rana's lab and launched it onto a 50m fiber. I got a few mW onto the fiber, enough to see with a card before disabling the laser. The fiber now runs along the hallway and terminates in rm 056. Its taped down everywhere someone might trip on it, but don't go out of your way to trip on it or pull on it because you are curious. Tomorrow I will co-run a BNC cable and attenuate the NPRO output so it can only send a few mW and so be laser safe. Then we can try to develop a procedure to align the beam to a suspended OMC and lock our suspended cavity goodness.

Notes to self: items needed from the 40m
  • ND10 and ND20 neutral density filter
  • EOM and mount set for 4 inch beam height
  • Post for fiber launch to get to 4 inch
  • Mode matching lens at 4in
  • 3x steering mirror at 4in
  • RF photodiode at 4in
  • Post for camera to 4in
  • Light sheild for camera
  • Long BNC cable
Some of these exist at 056 already
ELOG V3.1.3-